[figure: 9.7 cm]

NEW COPPER CHALCOGENIDES WITH MIXED VALENCE

I. Jacyna-Onyszkiewicz, M. Sidowski and S. Robaszkiewicz Institute of Physics, Adam Mickiewicz University, ul. Umultowska 85, 61-614 Poznań

We perform a comparative study of several new ternary copper chalcogenides with alkali metals and rare earth elements, synthesized recently. We focus on their electronic, magnetic and transport properties as well as the lattice structures. It turns out that in majority of these systems it is not copper but chalcogen which is mixed valent and the charge transport as well as charge orderings are dominated by the p-ligand holes. The most striking and unique features of these materials are: 1) The unusual linear in T behavior of metallic resistivity up to T ≪ Θ_{D} found in K_{3}Cu_{8}Te_{6}, KCu_{7−x}S_{4}, La_{2}CuS_{4−δ} and Y_{2}Cu_{2}S_{5−δ}, being similar to that observed in high-T_{c} oxides. 2) The occurrence of untypical metal-nonmetal transitions at T_{M} - in K_{3}Cu_{8}Se_{6} (T_{M} ∼ 160^{o}K), BaCuS_{3−x} (∼ 270^{o}K), K_{3}Cu_{8}Se_{6} (∼ 100^{o}K), TlCu_{7−x}S_{4} (∼ 180^{o}K), KCu_{7−x}S_{4} (∼ 190^{o}K) with, surprisingly, low temperature phases being metallic, while the high-temperature ones– semiconducting. 3) Phase transitions connected with CDW formation - in K_{3}Cu_{8}S_{6}, RbCu_{8}S_{6}, and K_{3}Cu_{8}Se_{6}. Electronic structure of the systems studied is discussed as a function of the charge transfer energy, the d-d Coulomb repulsion and the anion bandwidth.

[figure: 13.4 cm]

Subject category :

3. Transition Metals, Alloys and Compounds

Presentation mode : poster

Corresponding author : M. Sidowski

Address for correspondence : Institute of Physics, Adam Mickiewicz University, ul. Umultowska 85, 61-614 Poznań

Email address : masi@amu.edu.pl